OpenStudy (mathhelp346):

The distance that light travels in one year, 5.87*10^12, is called a light-year. If the distance from Proxima Centauri to Earth is 2.14*10^13 miles, how many years does it take light to travel from Proxima Centauri to Earth?

OpenStudy (dumbcow):

2.14 X10^13 = 21.4 x10^12 --> 21.4 / 5.87 = 3.65 years

OpenStudy (anonymous):

From my calcs 3.65 years is incorrect The speed of light is 3 x 10^8 m/s (3 x 10^8 m/s) x 60(sec) x 60(mins) x 24(hrs) x 365(days) = 9.46 x 10^15 meters/year ie the distance light travels in one year Proxima Centauri is 4 x 10^16 meters from earth Divide the distance to the star by the distance light travels in one year to give the amount of years that it takes light to travel from there to here (4 x 10^16) / (9.46 x 10^15) = 4.2 years for light to travel from that star Sorry in europe we use the SI system (why ye don't in America is incomprehensible to me), but use the same formula if needed for an assignment. After inspection and a little surfing, I think your figure for the distance to proxima centauri is closer to 2.5 x 10^13 miles. I don't know how accurate you need to be and I'm open to correction, but I think my figures can be verified through the internet at least if not through texts. PS the internet is brilliant, if you have not already done so you could have had the answer yourself in a matter of minutes. Instead of waiting for college students like me, who are trying to do anything to avoid studying, to do it for you :-) Glad to help all the same.